Basic Salary of an employee is Rs. 18,000 and Allowances are Rs. 9000. According to the company’s policy casual leaves are 12 days per year and earned leaves are 24 days per year and normal working days are 22 per month.
Find the total cost of leaves as percent of gross salary.

Answer by dkppathak(439)

Basic Salary of an employee is Rs. 18,000 and Allowances are Rs. 9000. According to the company’s policy casual leaves are 12 days per year and earned leaves are 24 days per year and normal working days are 22 per month.
Find the total cost of leaves as percent of gross salary.
total salary per month =18000+9000=27000
normal working days per months are 22x12=264 days and leave =12+24=36 per year
leave =36/264=9/66=3/22
total cost of leave=3x27000x12/22=972000/22=44181.81
percentage =44181.81x100/324000= 13.63%
leave salary 44181.81 and 13.63%